# Restoring Calibration for Aligned Large Language Models: A Calibration-Aware Fine-Tuning Approach

### v2

@@ 3 Preliminaries
Consider a probabilistic classification model $\boldsymbol{\hat{p}}:\mathcal{X}\rightarrow\Delta_{k}$ that outputs class probabilities for $k$ classes $1,\ldots,k$. For any given instance $x\in\mathcal{X}$ it would output some probability vector $\boldsymbol{\hat{p}}=(\hat{p}_{1}(x),\ldots,\hat{p}_{k}(x))$ belonging to $\Delta_{k}=\{(q_{1},\ldots,q_{k})\in[0,1]^{k}\mid\sum_{j=1}^{k}q_{j}=1\}$ which is the $(k-1)$-dimensional probability simplex over $k$ classes.
@@ 4.2 Calibratable and Non-Calibratable Regime
Figure 2: Illustration of TCE, calibratable and non-calibratable Regimes. Assume that the target probabilistic generative model $\boldsymbol{p}^{*}$ has a 60% accuracy. Left: The TCE range (yellow) is bounded between $C|60\%-\textnormal{Accuracy}|$ and $2\cdot|60\%-\textnormal{Accuracy}|$. Right: The calibratable regime (red) spans from 0 to TCE when accuracy $\leq 60\%$. The non-calibratable regime (blue) spans from a non-zero lower bound to TCE when accuracy $>60\%$.
@@ 5.2 Approximating the ECE loss for LLMs
For all $\boldsymbol{q}\in\Delta^{k}$, data points with $\boldsymbol{p}(x)=\boldsymbol{q}$ form a subset sharing the same label generation distribution. Since examining all values of $\boldsymbol{q}\in\Delta^{k}$ would be impractical, we make the following simplification. First, for all $\boldsymbol{q}\in\Delta^{k}$, we define $q=\arg\max\boldsymbol{q}$ and consider the subset $\{x|\max\boldsymbol{p}(x)=q\}$. Second, we discretize $q\in[0,1]$ into $M$ equal-width bins: $[0,1/M],\ldots,[(M-1)/M,1]$, consistent with traditional ECE estimation methods. Within the $m$th bin, the generative model assigns answers according to an unknown probability simplex where the largest probability equals $q_{m}$. Let $I=\arg\max\boldsymbol{p}(x)$. Then, we set $p_{j=I}(x)=q_{m}$. The probabilities $p_{j\neq I}(x)$ may vary across $x$ within the same bin. We apply a mapping to align these probabilities with the model’s confidence scores while maintaining the simplex constraints. We define $\boldsymbol{p}(x)$ as:
@@ 6.1 Experimental Setup
Table 2: Performance comparison among DPO/RLHF, Temperature Scaling, Label Smoothing, CFT, and RCFT across four models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in in-domain and out-domain scenarios. Best results in each metric block are bold. Blue highlights indicate superior in-domain conf-ECE of our CFT while red highlights denote best in-domain accuracy of our RCFT. $\downarrow$”/$\uparrow$” means the smaller/larger the better. “-” means the results of Temp. Scale. are the same as the original DPO/RLHF version.
Model Method conf-ECE $\downarrow$ cw-ECE $\downarrow$ Accuracy $\uparrow$
@@ A.3 Multiclass Calibration
[…]
@@ A.4 Rank Preserving Mapping
and since $\text{and since }\alpha=\frac{1-q_{m}}{T+3},\quad$ $\Rightarrow\quad\frac{1-q_{m}}{T+3}<\frac{q_{m}}{2}.$ (15)
@@ Appendix B Proof of Technical Results
[…]
@@ C.1 Comprehensive Descriptions for Models, Baselines, Datasets, and Metrics
[…]
@@
[…]
@@
Win Rate=Number of WinsTotal Number of $\text{Win Rate}=\frac{\text{Number of Wins}}{\text{Total Number of Comparisons}}.$
@@ C.4 Discussion of Bin Size
Model Method conf-ECE $\downarrow$ cw-ECE $\downarrow$

### v3

@@ 3 Preliminaries
Consider a probabilistic classification model $\boldsymbol{\hat{p}}:\mathcal{X}\rightarrow\Delta_{k}$ that outputs class probabilities for $k$ classes $1,\ldots,k$. For any given instance $x\in\mathcal{X}$ it would output some probability vector $\boldsymbol{\hat{p}}=(\hat{p}_{1}(x),\ldots,\hat{p}_{k}(x))$ belonging to $\Delta_{k}=\{(q_{1},\ldots,q_{k})\in[0,1]^{k}\mid\sum_{j=1}^{k}q_{j}=1\}$ which is the (k1)(k-1)-dimensional probability simplex over $k$ classes.
@@ 4.2 Calibratable and Non-Calibratable Regime
Figure 2: Illustration of TCE, calibratable and non-calibratable Regimes. Assume that the target probabilistic generative model $\boldsymbol{p}^{*}$ has a 60% accuracy. Left: The TCE range (yellow) is bounded between $C|60\%-\textnormal{Accuracy}|$ and $2\cdot|60\%-\textnormal{Accuracy}|$. Right: The calibratable regime (red) spans from 0 to TCE when accuracy $\leq 60\%$. The non-calibratable regime (blue) spans from a non-zero lower bound to TCE when accuracy $60\%$.
@@ 5.2 Approximating the ECE loss for LLMs
For all $\boldsymbol{q}\in\Delta^{k}$, data points with $\boldsymbol{p}(x)=\boldsymbol{q}$ form a subset sharing the same label generation distribution. Since examining all values of $\boldsymbol{q}\in\Delta^{k}$ would be impractical, we make the following simplification. First, for all $\boldsymbol{q}\in\Delta^{k}$, we define $q=\arg\max\boldsymbol{q}$ and consider the subset $\{x|\max\boldsymbol{p}(x)=q\}$. Second, we discretize $q\in[0,1]$ into $M$ equal-width bins: $0,1/M],\ldots,[(M-1)/M,1]$, consistent with traditional ECE estimation methods. Within the $m$th bin, the generative model assigns answers according to an unknown probability simplex where the largest probability equals $q_{m}$. Let $I=\arg\max\boldsymbol{p}(x)$. Then, we set $p_{j=I}(x)=q_{m}$. The probabilities $p_{j\neq I}(x)$ may vary across $x$ within the same bin. We apply a mapping to align these probabilities with the model’s confidence scores while maintaining the simplex constraints. We define $\boldsymbol{p}(x)$ as:
@@ 6.1 Experimental Setup
Table 2: Performance comparison among DPO/RLHF, Temperature Scaling, Label Smoothing, CFT, and RCFT across four models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in in-domain and out-domain scenarios. Best results in each metric block are bold. Blue highlights indicate superior in-domain conf-ECE of our CFT while red highlights denote best in-domain accuracy of our RCFT. $\downarrow$$\uparrow$” means the smaller/larger the better. “-” means the results of Temp. Scale. are the same as the original DPO/RLHF version.
Model
Method
conf-ECE $\downarrow$ cw-ECE $\downarrow$ Accuracy $\uparrow$
@@ Acknowledgments
We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017811.
@@ A.3 Multiclass Calibration
$\textnormal{mc-ECE}=\mathbb{E}_{\boldsymbol{\hat{p}}(x)}\frac{1}{k}\sum_{j=1}^{k}\big|\mathbb{P}(y=j|\boldsymbol{\hat{p}}(x))-\hat{p}_{j}(x)\big|.$
@@ A.4 Rank Preserving Mapping
$\text{and since }\alpha=\frac{1-q_{m}}{T+3},\quad$ $\Rightarrow\quad\frac{1-q_{m}}{T+3}<\frac{q_{m}}{2}.$ (15)
@@ Appendix B Proof of Technical Results
$\textnormal{mc-ECE}(\boldsymbol{p})=\mathbb{E}_{x}\frac{1}{k}\sum_{j=1}^{k}\big|\mathbb{P}(y=j|\boldsymbol{p}(x))-p_{j}(x)\big|=0.$
@@ C.1 Comprehensive Descriptions for Models, Baselines, Datasets, and Metrics
$\text{conf-ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}|\mathbb{P}(Y=I|\mathbf{x}\in B_{m})-\mathbb{E}[\pi_{\theta}(\mathbf{y}=I|\mathbf{x})|\mathbf{x}\in B_{m}]|,$ (19)
@@
$\text{cw-ECE}=\frac{1}{K}\sum_{j=1}^{k}\sum_{m=1}^{M}\frac{|B_{m,j}|}{N}|\mathbb{P}(Y=i|\mathbf{x}\in B_{m,j})-\mathbb{E}[\pi_{\theta}(\mathbf{y}=i|\mathbf{x})|\mathbf{x}\in B_{m,j}]|,$ (20)
@@
$\text{Win Rate}=\frac{\text{Number of Wins}}{\text{Total Number of Comparisons}}.$
@@ C.4 Discussion of Bin Size
Model
Method
conf-ECE $\downarrow$ cw-ECE $\downarrow$